It takes a shark 3 Seconds to accelerate at a rate of 4.3m/s2 To reach a final speed of 15.0 What was the sharks initial speed?
1
Expert's answer
2011-01-27T20:19:40-0500
Use the equation of the motion:
V = V0 +at
V0 = V-at = 15 - 3*4.3 = 2.1 m/s